# NeCo[Uncaptioned image]: Improving DINOv2’s spatial representations in 19 GPU hours with Patch Neighbor Consistency

### v1

@@ Abstract
We propose sorting patch representations across views as a novel self-supervised learning signal to improve pretrained representations. To this end, we introduce NeCo: Patch Neighbor Consistency, a novel training loss that enforces patch-level nearest neighbor consistency across a student and teacher model, relative to reference batches. Our method leverages a differentiable sorting method applied on top of pretrained representations, such as DINOv2-registers to bootstrap the learning signal and further improve upon them. This dense post-pretraining leads to superior performance across various models and datasets, despite requiring only 19 hours on a single GPU. We demonstrate that this method generates high-quality dense feature encoders and establish several new state-of-the-art results: +5.5 % and + 6% for non-parametric in-context semantic segmentation on ADE20k and Pascal VOC, and +7.2% and +5.7% for linear segmentation evaluations on COCO-Things and -Stuff.
@@ 1 Introduction
Dense self-supervised learning trains feature extractors to produce representations for every pixel or patch of an image without supervision. This field has seen substantial advancements in recent years, notably improving unsupervised semantic segmentation [1, 2, 3, 4, 5], object-centric representation learning [6], and other dense downstream tasks like object tracking and object detection [7, 8, 9, 2].
@@
One particularly interesting use-case of densely pretrained encoders was developed by Balazevic et. al. [10]. They propose to solve semantic segmentation by posing it as a nearest-neighbor retrieval problem utilizing the features of the spatial patches. This non-parametric method not only mirrors in-context learning in large language models (LLMs) [11] but also delivers rapid and robust performance, especially with limited data.
Inspired by this idea, we turn this evaluation approach on its head and propose to use nearest-neighbor retrieval as a training mechanism for encoders. By integrating nearest-neighbor relationships into the training phase, we aim to train models that can naturally cluster and distinguish intricate visual similarities. This method promises to yield models with deeply semantic spatial features specifically tailored for in-context tasks, enhancing their adaptability and robustness. This approach, while promising, presents two main challenges.
@@
The first is the source of supervision. In the case of evaluation, ground-truth labels are used, yet we are interested in obtaining better self-supervised representations. While previous works [10, 9] address this by essentially converting dense learning to image-level learning via learnable pooling of patches, we offer a more practical and versatile solution. We simply start from already image-level pretrained models and adapt them further. We term this stage dense post-pretraining and demonstrate that it is an effective and fast solution to this problem, taking only 19 hours on a single GPU for tuning for a ViT-S/16 model.
@@
The second challenge is the discrete nature of nearest-neighbor retrieval, which does not yield gradients. To overcome this, we apply a differentiable sorting method [12], originally proposed for ranking supervision, that we can use to backpropagate gradients. As we demonstrate empirically, this results in a more efficient and effective algorithm.
@@
Our method enforces Patch Neighbor Consistency, so we term it NeCo. We show that it can be applied on top of image-level pretrained models such as DINO [13], and densely trained ones like iBOT [14], Leopart [1], CrIBO [9] and DINOv2 [15, 16] to obtain superior features for in-context scene understanding. Despite not being as close to our NeCo  training task, our method also consistently improves on downstream benchmarks such as unsupervised semantic and linear segmentation.
@@
  • We demonstrate NeCo’s utility by applying it to six different backbones and evaluating it on four datasets and four evaluation protocols, achieving performance gains from 6% to 16%.
  • We set several new state-of-the-art performances, for example on the in-context segmentation benchmark of [10], we outperform the previous methods such as CrIbo and DINOv2 on Pascal VOC and ADE20k by 4% to 13% across different metrics.
@@ 2 Related Works
Dense Self-supervised Learning.
Dense self-supervised learning methods aim to generate categorizable representations at the pixel or patch level, rather than the image level. This field has gained significant attention [2, 1, 3, 4, 9, 10, 17, 18, 8, 19, 8, 7] due to the observation that image-level self-supervised methods [20, 21, 22, 23, 24, 25, 26, 27, 28, 15, 13, 29, 30, 31, 32, 33, 34] do not necessarily produce expressive dense representations [10, 8, 35, 36].
CroC [4] is a recent method that has proposed a dense self-supervised loss to address the issue. It applies joint clustering between different views, ensuring that cluster centers capturing the same object are similar. Leopart [1] improves dense representations by applying a dense clustering loss to pretrained models. Extending this concept, Timetuning [2] has demonstrated that finetuning pretrained backbones over the temporal dimension of unlabeled videos enhances dense reasoning capabilities. Recently, Hummingbird [10] has proposed a dense loss that leverages attention within and across images during training, showing strong in-context scene understanding during evaluation. CrIBo [9] takes this further by explicitly enforcing cross-image nearest neighbor consistency between image objects, achieving state-of-the-art results.
@@
Unsupervised Object Segmentation.
Several works specifically target unsupervised object segmentation [37, 38, 39, 40, 6, 41, 42, 43, 44]. The goal of these works is not to learn semantic patch-level representations; instead, they often utilize the existing information in frozen pretrained backbones and train another model to explicitly solve semantic segmentation. For instance, Seitzer et al. [37] train a slot-attention encoder and decoder module [45] to reconstruct DINO [13] pretrained features with a few slots for each input image. This process enables the creation of per-image object cluster maps, where each slot represents a distinct object or part of an object within the image, based on the features it predicts.
@@ 3 Patch Neighbor Consistency
Figure 1: NeCo  overview. Given an input image $I$, two augmentations $\tau_{1}$ and $\tau_{2}$ are applied to create two different views, which are processed by the teacher and student encoders, $\phi_{t}$ and $\phi_{s}$ respectively. The teacher encoder is updated using Exponential Moving Average (EMA). The encoded features are then aligned using ROI Align and compared with reference features $F_{r}$ obtained by applying $\phi_{s}$ to other batch images. Next, pairwise distances $D_{ij}$ between $F_{s}$ and $F_{r}$, as well as between $F_{t}$ and $F_{r}$, are computed using cosine similarity. These distances are then sorted using differentiable sorting and utilized to force nearest order consistency across the views through the NeCo loss.
The goal is to develop a feature space in which, for a given input, patches representing the same object exhibit similar features, whereas patches representing different objects show distinct features. A key challenge for a self-supervised method in this process is defining the similarity between image patches. Although patches from the same object (e.g. a cat) are expected to be more similar to each other than to those from different objects (e.g. a dog), they may still depict different parts of the object (e.g. a cat’s tail and legs). Both part- and whole-object representations have important applications. Therefore, to generate a feature space that can reliably represent both, merely establishing a single clustering per object is not enough. Instead, this requires accurately ordering similarities, such that patches representing the same object and its parts are more similar to one another and less to those of other objects. To this end, our method works by extracting dense features of the inputs, finding their pair-wise distances, and forcing a consistency between the order of nearest neighbors within a batch across two views. Figure 1 provides an overview of the method, which we describe in detail below.
@@
We utilize the Vision Transformer (ViT) architecture [46] as the backbone and employ a teacher-student framework, where the student and teacher models are denoted by $\phi_{s}$ and $\phi_{t}$, respectively. The teacher’s weights are updated using the exponential moving average of the student’s weights.
As the generated views cover different parts of the input, the extracted features do not necessarily correspond to the same objects. To address this, we align the features by applying ROI-Align [47], adjusted according to the crop augmentation parameters. This process creates spatially aligned dense features for the teacher and student networks, represented by $F_{s}\in\mathbb{R}^{N^{\prime}\times d}$ and $F_{t}\in\mathbb{R}^{N^{\prime}\times d}$. These features are then forced to maintain a consistent order of nearest neighbors, ensuring more robust and meaningful feature representations, as explained next.
@@
To identify the nearest neighbors of the patches, it is necessary to extract features from other images in the batch and compute their distances with respect to $F_{s}$ and $F_{t}$. To achieve this, all batch images are fed through the student network, $\phi_{s}$, to obtain the batch features $F_{B}\in\mathbb{R}^{BN\times d}$. We sample a random fraction $f\ll 1$ of these patches to obtain the $R=fBN$ reference patches $F_{r}\in\mathbb{R}^{r\times d}$ which we use to compare the nearest neighbors of our $F_{s}$ and $F_{t}$ features. To this end, we compute distances based on cosine similarities,
@@
To determine the order of nearest neighbors from distance matrices, sorting is necessary. However, traditional sorting algorithms cannot propagate gradients because they use non-differentiable operations such as $d^{\prime}_{i}\leftarrow\min(d_{a},d_{b})$ and $d^{\prime}_{a}\leftarrow\max(d_{a},d_{b})$ to facilitate element swapping in the sequence for an ordering $a<b$. Given a sequence $s=\bigl{(}d_{1},\ldots,d_{R}\bigr{)}$, where $R$ is the length of the sequence, We use relaxed, differentiable versions of these operations by defining their soft versions following recent work [17], as follows:
@@
where the function $f(x)=\frac{1}{2}(\arctan(\beta x)+0.5)$, and $\beta>0$ is an inverse temperature parameter, specifying the steepness of the operator. This function is sigmoid-shaped and centered around $x=0$. As $\beta$ approaches infinity, the relaxation converges to the discrete swap operation. This operation can be defined in an approximate permutation matrix $P_{\text{swap}}(d_{i},d_{j})\in\mathbb{R}^{L\times L}$, which is essentially an identity matrix except for the entries $P_{ii}$, $P_{ij}$, $P_{ji}$, and $P_{jj}$ defined as
@@
where $M$ is the set of odd indices if $t$ is odd and the set of even indices if $t$ is even. The overall permutation matrix $Q$ is obtained by multiplying the permutation matrices from all steps of the sorting algorithm, $Q=\prod_{t=1}^{L}P_{t}.$ As shown by [12], $L=R$ of such steps are sufficient for efficient sorting. In the discrete case, for each column $i$, the permutation matrix has exactly one entry of 1, indicating the sequence element that should be placed in the $i$-th column. In the relaxed version, column values represent a distribution over possible sequence elements. In our case, a row $i$ of the distance matrix $D_{s}$ shows the distance of the $i$-th student feature to all the reference features.
@@ 4.1 Setup
Benchmarked Methods.
We compare our method against state-of-the-art dense self-supervised learning methods, including CrIBo [9], Hummingbird [10], TimeT [2], and Leopart [1]. To provide a more comprehensive evaluation, We also include the performance of DINOv2 enhanced with registers, referred to as DINOv2R [15, 16], as it has demonstrated strong dense capabilities. Additionally, we benchmark our method against leading unsupervised semantic segmentation approaches such as COMUS [44].
Training.
We run our experiments on ViT-Small and ViT-Base with a patch size of 14. We start from various pretrained backbones, and use DINOv2 with registers’ unless otherwise noted. We post-pretrain these models for 25 COCO epochs on a single NVIDIA RTX A6000-46GB GPU, taking around 19 hours. For other training details, we refer readers to the Appendix A.
Evaluation.
In all our evaluations, we discard the projection head, following previous works [13, 1, 2], and directly use the spatial tokens from the Vision Transformer backbone. Scores in all experiments are reported as mean intersection over union (mIoU). We conduct four types of evaluations: linear segmentation fine-tuning with a $1\times 1$ convolution, end-to-end segmentation with the Segmenter head [48], clustering and overclustering semantic segmentation [1, 2], and dense nearest neighbor retrieval [10]. For clustering and overclustering, we apply K-Means to spatial tokens, setting $K$ to the number of ground truth objects and to high values like 300 and 500, as previously used [1, 2]. We then extract object cluster maps and match them using Hungarian matching [49]. For dense nearest neighbor retrieval, we follow the protocol from [10], implemented in [50].
Datasets.
We train our model on ImageNet-100 [51], Pascal VOC12 [52], and COCO [53] for ablations and use COCO as our primary training dataset for all state-of-the-art comparisons. For evaluations, we use the validation sets of Pascal VOC12 [52], COCO [53], ADE20k [54], and Pascal Context [55]. For finetuning and feature transferability evaluations on COCO [56], we train using a 10% split of the training set, while we use the full training splits of the other datasets.

### v2

# Near, far: Patch-ordering enhances vision foundation models’ scene understanding

@@ Abstract
We introduce NeCo: Patch Neighbor Consistency, a novel self-supervised training loss that enforces patch-level nearest neighbor consistency across a student and teacher model. Compared to contrastive approaches that only yield binary learning signals, i.e. ‘attract’ and ‘repel’, this approach benefits from the more fine-grained learning signal of sorting spatially dense features relative to reference patches. Our method leverages differentiable sorting applied on top of pretrained representations, such as DINOv2-registers to bootstrap the learning signal and further improve upon them. This dense post-pretraining leads to superior performance across various models and datasets, despite requiring only 19 hours on a single GPU. This method generates high-quality dense feature encoders and establishes several new state-of-the-art results such as +5.5 % and +6% for non-parametric in-context semantic segmentation on ADE20k and Pascal VOC, +7.2% and +5.7% for linear segmentation evaluations on COCO-Things and -Stuff and improvements in the 3D understanding of multi-view consistency on SPair-71k, by more than 1.5%.
@@ 1 Introduction
Dense self-supervised learning trains feature extractors to produce representations for every pixel or patch of an image without supervision. This field has seen substantial advancements in recent years, notably improving unsupervised semantic segmentation (Ziegler & Asano, 2022; Salehi et al., 2023; Araslanov et al., 2021; Stegmüller et al., 2023; Wang et al., 2021), object-centric representation learning (Zadaianchuk et al., 2023), and other dense downstream tasks like object tracking and object detection (Hénaff et al., 2021; 2022; Lebailly et al., 2023; Salehi et al., 2023).
@@
One particularly interesting use-case of densely pretrained encoders was developed by Balazevic et al. (2023). They propose to solve semantic segmentation by posing it as a nearest-neighbor retrieval problem utilizing the features of the spatial patches. This non-parametric method not only mirrors in-context learning in large language models (LLMs) (Brown et al., 2020) but also delivers rapid and robust performance, especially with limited data.
Building on this idea, we propose an inverted approach: using nearest-neighbor retrieval not just for evaluation but as a training mechanism for encoders. This approach promises a more fine-grained learning signal, enabling the capture of intricate visual details. For instance, the features of a tire should be closely related to one another, as well as to those of a car body, while remaining distinct from features of an airplane.
Unlike contrastive losses, which offer a binary ‘attract’ or ‘repel’ signal, this provides a much richer, continuous learning signal. Additionally, it avoids the pitfalls of reconstruction-based methods like MAE (He et al., 2022), where low-level RGB patches that are visually similar may not carry similar semantics. By operating strictly in the deep feature space, learning is guided by higher-level semantics rather than superficial pixel-level similarities. As a result, this method promises to yield models with deeply semantic spatial features specifically tailored for in-context tasks, enhancing their adaptability and robustness. However, this approach, while promising, presents two main challenges.
@@
The first is the source of supervision. In the case of evaluation, ground-truth labels are used, yet we are interested in obtaining better self-supervised representations. While previous works (Balazevic et al., 2023; Lebailly et al., 2023) address this by essentially converting dense learning to image-level learning via learnable pooling of patches, we offer a more practical and versatile solution. We simply start from already image-level pretrained models and adapt them further. We term this stage dense post-pretraining and demonstrate that it is an effective and fast solution to this problem, taking only 19 hours on a single GPU for tuning for a ViT-S/16 model.
@@
The second challenge is the discrete nature of nearest-neighbor retrieval, which does not yield gradients. To overcome this, we apply a differentiable sorting method proposed by Petersen et al. (2021), originally developed for ranking supervision, that we can use to backpropagate gradients. As we demonstrate empirically, this results in a more efficient and effective algorithm.
@@
Our method enforces Patch Neighbor Consistency, so we term it NeCo. We show that it can be applied on top of image-level pretrained models such as DINO (Caron et al., 2021), and densely trained ones like iBOT (Zhou et al., 2022), Leopart (Ziegler & Asano, 2022), CrIBO (Lebailly et al., 2023) and DINOv2 (Oquab et al., 2023; Darcet et al., 2024) to obtain superior features for in-context scene understanding. Despite not being as close to our NeCo  training task, our method also consistently excels on downstream benchmarks such as 3D understanding, unsupervised semantic segmentation and also full-finetuning semantic segmentation, where it even improves upon the state-of-the-art DINOv2-R model (Darcet et al., 2024).
@@
  • We demonstrate NeCo’s utility by applying it to six different backbones and evaluating it on five datasets and five evaluation protocols, achieving performance gains from 6% to 16%.
  • We set several new state-of-the-art performances, for example on the in-context segmentation benchmark of Balazevic et al. (2023), we outperform the previous methods such as CrIbo and DINOv2 on Pascal VOC and ADE20k by 4% to 13% across different metrics.
@@ 2 Related Works
Dense Self-supervised Learning. Dense self-supervised learning methods aim to generate categorizable representations at the pixel or patch level, rather than the image level. This field has gained significant attention (Salehi et al., 2023; Ziegler & Asano, 2022; Araslanov et al., 2021; Stegmüller et al., 2023; Lebailly et al., 2023; Balazevic et al., 2023; Hwang et al., 2019; Liu et al., 2020; Hénaff et al., 2022; Van Gansbeke et al., 2021; Hénaff et al., 2022; 2021; Yun et al., 2022) due to the observation that image-level self-supervised methods (Caron et al., 2018; Asano et al., 2019; Chen et al., 2020; Grill et al., 2020; Caron et al., 2021; Izacard et al., 2021; Oquab et al., 2023) do not necessarily produce expressive dense representations (Balazevic et al., 2023; Hénaff et al., 2022; He et al., 2019; Purushwalkam & Gupta, 2020).
CroC (Stegmüller et al., 2023) is a recent method that has proposed a dense self-supervised loss to address the issue. It applies joint clustering between different views, ensuring that cluster centers capturing the same object are similar. Leopart (Ziegler & Asano, 2022) improves dense representations by applying a dense clustering loss to pretrained models. Extending this concept, TimeTuning (Salehi et al., 2023) has demonstrated that finetuning pretrained backbones over the temporal dimension of unlabeled videos enhances dense reasoning capabilities. Recently, Hummingbird (Balazevic et al., 2023) has proposed a dense loss that leverages attention within and across images during training, showing strong in-context scene understanding during evaluation. CrIBo (Lebailly et al., 2023) takes this further by explicitly enforcing cross-image nearest neighbor consistency between image objects, achieving state-of-the-art results.
@@
Unsupervised Object Segmentation. Several works specifically target unsupervised object segmentation (Seitzer et al., 2022; Löwe et al., 2023; Bao et al., 2023; Siméoni et al., 2023; Zadaianchuk et al., 2023; Wang et al., 2023; Hamilton et al., 2022; Lan et al., 2023; Zadaianchuk et al., 2022). The goal of these works is not to learn semantic patch-level representations; instead, they often utilize the existing information in frozen pretrained backbones and train another model to explicitly solve semantic segmentation. For instance, Seitzer et al. (2022) train a slot-attention encoder and decoder module (Locatello et al., 2020) to reconstruct DINO (Caron et al., 2021) pretrained features with a few slots for each input image. This process enables the creation of per-image object cluster maps, where each slot represents a distinct object or part of an object within the image, based on the features it predicts.
@@ 3 Patch Neighbor Consistency
Figure 1: NeCo  overview. Given an input image $I$, two augmentations $\tau_{1}$ and $\tau_{2}$ are applied to create two different views, which are processed by the teacher and student encoders, $\phi_{t}$ and $\phi_{s}$ respectively. The teacher encoder is updated using Exponential Moving Average (EMA). The encoded features are then aligned using ROI Align and compared with reference features $F_{r}$ obtained by applying $\phi_{t}$ to other batch images. Next, pairwise distances $D_{ij}$ between $F_{s}$ and $F_{r}$, as well as between $F_{t}$ and $F_{r}$, are computed using cosine similarity. These distances are then sorted using differentiable sorting and utilized to force nearest order consistency across the views through the NeCo loss.
The goal is to develop a feature space in which, for a given input, patches representing the same object exhibit similar features, whereas patches representing different objects show distinct features. A key challenge for a self-supervised method in this process is defining the similarity between image patches. Although patches from the same object (e.g. a cat) are expected to be more similar to each other than to those from different objects (e.g. a dog), they may still depict different parts of the object (e.g. a cat’s tail and legs).
Therefore, a learned dense feature space must provide an ordering of similarities to ensure that patches from the same object and its parts are correctly distinguished from those of other objects. To this end, our method works by extracting dense features of the inputs, finding their pair-wise distances, and forcing a consistency between the order of nearest neighbors within a batch across two views. While contrastive and clustering-based methods (Chen et al., 2020; Caron et al., 2020; Ziegler & Asano, 2022; Salehi et al., 2023) can be seen as potential substitutes for sorting in aligning positive views, they only enforce similarity without capturing the nuanced ordering constraints necessary for structured relationships. Moreover, they often rely on explicit negative samples, which limits their flexibility. In contrast, our sorting-based approach not only maintains relative similarities without the need for negatives but also prevents mode collapse by preserving subtle distinctions among patches, even within the same object. Figure 1 provides an overview of the method, which we describe in detail below.
@@
We utilize the Vision Transformer (ViT) architecture (Dosovitskiy et al., 2020) as the backbone and employ a teacher-student framework, where the student and teacher models are denoted by $\phi_{s}$ and $\phi_{t}$, respectively. The teacher’s weights are updated using the exponential moving average of the student’s weights.
As the generated views cover different parts of the input, the extracted features do not necessarily correspond to the same objects. To address this, we align the features by applying ROI-Align (He et al., 2017), adjusted according to the crop augmentation parameters. This process creates spatially aligned dense features for the teacher and student networks, represented by $F_{s}\in\mathbb{R}^{N^{\prime}\times d}$ and $F_{t}\in\mathbb{R}^{N^{\prime}\times d}$. These features are then forced to maintain a consistent order of nearest neighbors, ensuring more robust and meaningful feature representations, as explained next.
@@
To identify the nearest neighbors of the patches, it is necessary to extract features from other images in the batch and compute their distances with respect to $F_{s}$ and $F_{t}$. To achieve this, all batch images are fed through the teacher network, $\phi_{t}$, to obtain the batch features $F_{B}\in\mathbb{R}^{BN\times d}$. We sample a random fraction $f\ll 1$ of these patches to obtain the $R=fBN$ reference patches $F_{r}\in\mathbb{R}^{r\times d}$ which we use to compare the nearest neighbors of our $F_{s}$ and $F_{t}$ features. To this end, we compute distances based on cosine similarities,
@@
To determine the order of nearest neighbors from distance matrices, sorting is necessary. However, traditional sorting algorithms cannot propagate gradients because they use non-differentiable operations such as $d^{\prime}_{i}\leftarrow\min(d_{a},d_{b})$ and $d^{\prime}_{a}\leftarrow\max(d_{a},d_{b})$ to facilitate element swapping in the sequence for an ordering $a<b$. Given a sequence $s=\bigl{(}d_{1},\ldots,d_{R}\bigr{)}$, where $R$ is the length of the sequence, We use relaxed, differentiable versions of these operations by defining their soft versions following recent work (Lee et al., 2017), as follows:
@@
where the function $f(x)=\frac{1}{\pi}\arctan(\beta x)+0.5$, and $\beta>0$ is an inverse temperature parameter, specifying the steepness of the operator. This function is sigmoid-shaped and centered around $x=0$. As $\beta$ approaches infinity, the relaxation converges to the discrete swap operation. This operation can be defined in an approximate permutation matrix $P_{\text{swap}}(d_{i},d_{j})\in\mathbb{R}^{L\times L}$, which is essentially an identity matrix except for the entries $P_{ii}$, $P_{ij}$, $P_{ji}$, and $P_{jj}$ defined as
@@
where $M$ is the set of odd indices if $t$ is odd and the set of even indices if $t$ is even. The overall permutation matrix $Q$ is obtained by multiplying the permutation matrices from all steps of the sorting algorithm, $Q=\prod_{t=1}^{L}P_{t}.$ As shown by (Petersen et al., 2021), $L=R$ of such steps are sufficient for efficient sorting. In the discrete case, for each column $i$, the permutation matrix has exactly one entry of 1, indicating the sequence element that should be placed in the $i$-th column. In the relaxed version, column values represent a distribution over possible sequence elements. In our case, a row $i$ of the distance matrix $D_{s}$ shows the distance of the $i$-th student feature to all the reference features.
@@ 4.1 Setup
Benchmarked Methods. We compare our method against state-of-the-art dense self-supervised learning methods, including CrIBo (Lebailly et al., 2023), Hummingbird (Balazevic et al., 2023), TimeT (Salehi et al., 2023), Leopart (Ziegler & Asano, 2022), and CrOC (Stegmüller et al., 2023) as well as baselines such ad DINO (Caron et al., 2021), iBOT (Zhou et al., 2022). To provide a more comprehensive evaluation, We also include the performance of DINOv2 enhanced with registers, referred to as DINOv2R (Oquab et al., 2023; Darcet et al., 2024), as it has demonstrated strong dense capabilities. Additionally, we benchmark our method against leading unsupervised semantic segmentation approaches such as COMUS (Zadaianchuk et al., 2022), DINOSAUR (Seitzer et al., 2022), DeepSpectral (Melas-Kyriazi et al., 2022), and MaskContrast (Van Gansbeke et al., 2021).
Training. We run our experiments on ViT-Small and ViT-Base with a patch size of 14. We start from various pretrained backbones, and use DINOv2 with registers unless otherwise noted. We post-pretrain these models for 25 COCO epochs on a single NVIDIA RTX A6000-46GB GPU, taking around 19 hours. For other training details, including detailed computational efficiency analysis, we refer readers to the Appendix A.
Evaluation. In all our evaluations, we discard the projection head, following previous works (Caron et al., 2021; Ziegler & Asano, 2022; Salehi et al., 2023), and directly use the spatial tokens from the Vision Transformer backbone. Scores in all experiments are reported as mean intersection over union (mIoU). We conduct four types of evaluations: linear segmentation fine-tuning with a $1\times 1$ convolution, end-to-end segmentation with the Segmenter head (Strudel et al., 2021), clustering and overclustering semantic segmentation (Ziegler & Asano, 2022; Salehi et al., 2023), and dense nearest neighbor retrieval (Balazevic et al., 2023). For clustering and overclustering, we apply K-Means to spatial tokens, setting $K$ to the number of ground truth objects and to high values like 300 and 500, as previously used (Ziegler & Asano, 2022; Salehi et al., 2023). We then extract object cluster maps and match them using Hungarian matching (Kuhn, 1955). For dense nearest neighbor retrieval, we follow the protocol from Balazevic et al. (2023), implemented in Pariza et al. (2024). For 3D understanding benchmark, We use the multiview feature consistency evaluation method proposed by El Banani et al. (2024), except all the images are resized to 224 $\times$ 224.
Datasets. We train our model on ImageNet-100 (Tian et al., 2020), Pascal VOC12 (Everingham et al., 2010), and COCO (Lin et al., 2014) for ablations and use COCO as our primary training dataset for all state-of-the-art comparisons. For evaluations, we use the validation sets of Pascal VOC12 (Everingham et al., 2010), COCO (Lin et al., 2014), ADE20k (Zhou et al., 2017), and Pascal Context (Mottaghi et al., 2014). For finetuning and feature transferability evaluations on COCO (Caesar et al., 2018), we train using a 10% split of the training set, while we use the full training splits of the other datasets. For the 3D understanding benchmark, we use Spair-71k (Min et al., 2019) dataset.